OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -10 - 2 + 6 + 14 + ... + 110

OpenStudy (perl):

keep adding 8

OpenStudy (lucaz):

first you need to know the number of terms, 110 is what term in the sequence: \[\Large{ a_n=a_1+d(n-1)\\ 110=-10+8(n-1)\\ 110=-18+8n\\ n=\frac{128}{8}=16 }\]

OpenStudy (perl):

the terms of your sum is an arithmetic sequence an = a1 + (n-1)*d an = -10 + (n-1)*8 an = -10 + 8n - 8 an = 8n -18

OpenStudy (perl):